If the letters of the word MISSISSIPPI are written down at random in a row, the probability that no two S's come together is?

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
In this arrangement we can use gap method.
 
(7! /4! 2! ) * 8C4
 
 = (7 *6 *5 /2) * 8*7*6*5 /24
 
 = 105 * 70 = 7350
 
 
Regards
Arun